# Quantum and Classical Communication Complexity of Permutation-Invariant Functions

### v1

@@ 1.3 Related Work
In previous work, Ghazi, Kamath and Sudan [24] introduced a complexity measure, which is polynomially equivalent to the randomized communication complexity of permutation-invariant functions defined in Definition 2.1. This paper is inspired by their work.
@@ 2 Preliminaries
[…]
@@ 2.1 Boolean Functions
A partial function is a function defined only on a subset of its domain $\mathcal{X}$. Formally, given a partial Boolean function $f\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$, $f(x)$ is undefined for $x\in\mathcal{X}$ if $f(x)=*$. A total function is a function that is defined on the entire domain. We say $f\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$ is a subfunction of $g\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$ if $f(x)=g(x)$ or $f(x)=*$ for any $x\in\mathcal{X}$.
A Boolean predicate is a partial function that has domain $\mathcal{X}=\left\{0,1,...,n\right\}$ for any $n\in\mathbb{N}$.
An incomplete Boolean matrix is a matrix with entries in $\left\{-1,1,*\right\}$, where undefined entries are filled with ***.
@@
  • $\overline{x}\coloneqq\left\{\overline{x_{0}},...,\overline{x_{n-1}}\right\}=\left\{1-x_{0},...,1-x_{n-1}\right\}$;
  • $x\land y\coloneqq\left\{x_{0}\land y_{0},...,x_{n-1}\land y_{n-1}\right\}$; and
  • $x\oplus y\coloneqq\left\{x_{0}\oplus y_{0},...,x_{n-1}\oplus y_{n-1}\right\}$.
@@ 2.2 Communication Complexity Model
In the two-party communication model, Alice is given input $x$, and Bob is given input $y$. Then they aim to compute $f(x,y)$ for some function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ by communication protocols. The deterministic communication complexity $D(f)$ is defined as the cost of the deterministic protocol with the smallest communication cost, which computes $f$ correctly on any input. The randomized communication complexity $R_{\epsilon}(f)$ is defined as the cost of the randomized protocol with the smallest communication cost, which has access to public randomness and computes $f$ correctly on any input with probability at least $1-\epsilon$. Similarly, the quantum communication complexity $Q(f)$ is defined as the cost of the quantum protocol with the smallest cost, which is not allowed to share prior entanglement, has access to public randomness and computes $f$ correctly on any input with probability at least $1-\epsilon$. If the quantum protocol is allowed with prior entanglement initially, then the corresponding quantum communication complexity is denoted $Q^{*}(f)$. If a protocol succeeds with probability at least $1-\epsilon$ on any input for some constant $\epsilon<1/2$, we say the protocol is with bounded error. If $\epsilon=1/3$, we abbreviate $R_{\epsilon}(f),Q_{\epsilon}(f),Q_{\epsilon}^{*}(f)$ as $R(f),Q(f),Q^{*}(f)$.
@@ 2.3 Permutation-Invariant Functions
In the two-party communication model, the function value of a permutation-invariant function is invariant if we perform the same permutation to the inputs of Alice and Bob. Specifically, the formal definition is as follows.
Definition 2.1 (Permutation-invariant functions [24]).
Note that any permutation-invariant function $f$ in Definition 2.1 depends only on $|x|,|y|$ and $|x\land y|$. Here ||$\cdot$|| ⋅ | is the Hamming weight for the binary string, i.e., the number of 1’s in the string. Thus, for any $a,b\in[n]$, there exists a function $f_{a,b}\colon\left\{\max\left\{0,a+b-n\right\},...,\min\left\{a,b\right\}\right\}\rightarrow\left\{-1,1\right\}$ such that
@@
Definition 2.2 (Jump in $f_{a,b}$ [24]).
$(c,g)$ is a jump in $f_{a,b}$ if
@@
[…]
The following definition gives an important instance of permutation-invariant functions.
Definition 2.3 (Set-Inclusion Problem).
We define the Set-Inclusion Problem $\mathsf{SetInc}^{n}_{a,b,c,g}$ as the following partial function:
$\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $|x\land y|\leq c-g$},\\ 1&\text{if $|x|=a,|y|=b$ and $|x\land y|\geq c+g$},\\ *&\text{otherwise.}\end{cases}$
Additionally, we define the Exact Set-Inclusion Problem $\mathsf{ESetInc}_{a,b,c,g}^{n}$ as follows.
$\mathsf{ESetInc}_{a,b,c,g}^{n}(x,y)\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $|x\land y|=c-g$},\\ 1&\text{if $|x|=a,|y|=b$ and $|x\land y|=c+g$},\\ *&\text{otherwise.}\end{cases}$
2.4 Rank and Approximate Rank
If $F$ is a real matrix, let $\operatorname{\mathrm{rank}}(F)$ be the rank of $F$. Then we define the approximate rank for any incomplete matrix as follows.
Definition 2.4 (Approximate rank).
For an incomplete matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1, we say a real matrix $A$ approximates $F$ with error $\epsilon$ if:
  1. (1)
    $|A_{i,j}-F_{i,j}|\leq\epsilon$ for any $i\in[m],j\in[n]$ such that $F_{i,j}\neq*$;
  2. (2)
    $|A_{i,j}|\leq 1$ for all $i\in[m],j\in[n]$.
Let $\mathcal{F}_{\epsilon}$ be the set of all the real matrices that approximate $F$ with error $\epsilon$. The approximate rank of $F$ with error $\epsilon$, denoted by $\operatorname{\widetilde{rank}}_{\epsilon}(F)$, is the least rank among all real matrices in $\mathcal{F}_{\epsilon}$. If $\epsilon=2/3$, we abbreviate $\operatorname{\widetilde{rank}}_{\epsilon}(F)$ as $\operatorname{\widetilde{rank}}(F)$.
Let $f$ be a Boolean function, $\operatorname{\mathrm{rank}}(f)\coloneqq\operatorname{\mathrm{rank}}\left(M_{f}\right)$ and $\operatorname{\widetilde{rank}}(f)\coloneqq\operatorname{\widetilde{rank}}{\left(M_{f}\right)}$, where $M_{f}$ is the input matrix of $f$.
@@ 3 Polynomial Equivalence on Communication Complexity of Permutation-Invariant Functions
To show the polynomial equivalence between quantum and classical communication complexity of permutation-invariant functions as stated in Theorem 1.5, we prove the following two theorems (proved in Sections 3.1 and 3.2, respectively) for the quantum and randomized communication complexities of permutation-invariant functions using the measure in Definition 3.1.
Definition 3.1 (Measure $m(f)$).
Fix $n\in\mathbb{Z}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1,*\}$ be a permutation-invariant function. We define the measure $m(f)$ of $f$ as follows:
$m(f)\coloneqq\max_{\begin{subarray}{c}a,b\in[n]\\ (c,g)\in\mathscr{J}\left(f_{a,b}\right)\\ n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}\\ n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)\end{subarray}}\frac{\sqrt{n_{1}n_{2}}}{g}\enspace.$
Note that this definition is motivated by Lemma 3.4.
The measure $m(\cdot)$ is inspired by the complexity measure introduced in [24], which was used to capture the randomized communication complexity of permutation-invariant functions.
@@
Fix $n\in\mathbb{Z}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1,*\}$ be a permutation-invariant function. We have
$Q^{*}(f)=\Omega(m(f))\enspace.$
@@
Fix $n\in\mathbb{N}$. Given a permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1,*\}$ and the corresponding measure $m(f)$ defined in Definition 3.1, we have
  • $R(f)=O\left(m(f)^{2}\log^{2}n\log\log n\right)$, and
  • $Q(f)=O\left(m(f)\log^{2}n\log\log n\right)$.

### v2

@@ 1.4 Comparison with previous work
In previous work, Ghazi, Kamath and Sudan [23] considered the randomized communication complexity of permutation-invariant functions. For any permutation-invariant function $f$, they introduced a complexity measure (denoted by $M(f)$) almost equivalent up to a fourth power of $R(f)$, i.e.,
$R(f)$ $=\Omega(M(f)),$
$R(f)$ $=O\left(M(f)^{4}\log M(f)\log\log n\log\log\log n+\log n\right).$
In this paper, we propose a new complexity measure $m(f)$, refining the argument in [23], which is almost quadratically equivalent to $R(f)$ and tightly characterizes $Q(f)$ up to a polylogarithmic factor of the input size (See Theorem 1.8). This enables us to prove a quadratic equivalence between $R(f)$ and $Q(f)$ for all permutation-invariant functions $f$.
Specifically, [23] first considered the Gap-Hamming-Distance Problem as Definition 1.12, and then generalized the result to all permutation-invariant functions. In our work, we use a similar framework. The difference point is that, instead of the Gap-Hamming-Distance Problem, we consider the Set-Inclusion Problem as Definition 1.13. It is worth noting that, while the two definition forms are different, the Set-Inclusion Problem is equivalent to the Gap-Hamming-Distance Problem, i.e., $\mathsf{SetInc}_{a,b,c,g}^{n}$ is equivalent to $\mathrm{GHD}_{a,b,a+b-2c,2g}^{n}$.
Definition 1.12 (Gap-Hamming-Distance Problem [23]).
Fix $n\in\mathbb{Z}^{+}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming distances of $\Delta(x,y)$ when $|x|=a,|y|=b$. The Gap-Hamming-Distance Problem $\mathrm{GHD}^{n}_{a,b,c,g}:\left\{0,1\right\}^{n}\times\left\{0,1\right\}^{n}\rightarrow\left\{-1,1,*\right\}$ is defined as the following partial function:
$\mathrm{GHD}_{a,b,c,g}^{n}(x,y)$
$\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $\Delta(x,y)\geq c+g$},\\ 1&\text{if $|x|=a,|y|=b$ and $\Delta(x,y)\leq c-g$},\\ *&\text{otherwise,}\end{cases}$
where $\Delta(x,y)=|\left\{i\in\{0,1,...,n-1\}|x_{i}\neq y_{i}\right\}|$ is the Hamming distance between $x$ and $y$. Additionally, the Exact Gap-Hamming-Distance Problem $\overline{\mathrm{GHD}}_{a,b,c,g}^{n}$ as follows.
$\overline{\mathrm{GHD}}_{a,b,c,g}^{n}(x,y)$
$\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $\Delta(x,y)=c-g$},\\ 1&\text{if $|x|=a,|y|=b$ and $\Delta(x,y)=c+g$},\\ *&\text{otherwise.}\end{cases}$
Definition 1.13 (Set-Inclusion Problem).
Fix $n\in\mathbb{Z}^{+}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming weights of $|x\land y|$ when $|x|=a,|y|=b$. The Set-Inclusion Problem $\mathrm{SetInc}^{n}_{a,b,c,g}:\left\{0,1\right\}^{n}\times\left\{0,1\right\}^{n}\rightarrow\left\{-1,1,*\right\}$ is defined as the following partial function:
$\mathrm{SetInc}_{a,b,c,g}^{n}(x,y)$
$\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $|x\land y|\leq c-g$},\\ 1&\text{if $|x|=a,|y|=b$ and $|x\land y|\geq c+g$},\\ *&\text{otherwise.}\end{cases}$
Additionally, the Exact Set-Inclusion Problem $\mathrm{ESetInc}_{a,b,c,g}^{n}$ is defined as follows.
$\mathrm{ESetInc}_{a,b,c,g}^{n}(x,y)$
$\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $|x\land y|=c+g$},\\ 1&\text{if $|x|=a,|y|=b$ and $|x\land y|=c-g$},\\ *&\text{otherwise.}\end{cases}$
Next, we show the difference between [23] and our result about the Set-Inclusion Problem (or Gap-Hamming-Distance Problem equivalently) from upper and lower bounds respectively.
For the upper bound, we designed different protocols from [23]. The details of our protocols are shown in 3.2. Moreover, the comparison of required communication costs is shown in Table 1.
For the lower bound, noting that the information complexity of the Unique-Set-Disjointness problem ($\mathrm{UDISJ}$) has been known [6], by a series of reductions, [23] showed the information complexity of UDisj is a lower bound on the randomized communication complexity of the Exact Gap-Hamming-Distance Problem (See Section 3.3 in [23]). As a comparison, we use a different reduction method as Figure 1.
$\mathrm{ESetInc}_{a,b,c,g}^{n}$$\mathrm{ESetInc}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}^{n_{1}+3n_{2}}$$\mathrm{ESetInc}_{2k,k,l,1/2}^{4k}$Lemma 3.4Lemma 3.5
Figure 1: Our reduction path for the lower bound of $\mathsf{ESetInc}$.
We first reduce the lower bound of the Exact Set-Inclusion Problem to some smaller instances with specific parameters (See Lemmas 3.4 and 3.5). Then we resolve the lower bound on the quantum communication complexity of the instance $\mathrm{ESetInc}_{2k,k,l,1/2}^{4k}$ (See Lemma 3.7). Compared to [23], our reduction is more concise and direct.
For the communication complexity of $\mathsf{SetInc}_{a,b,c,g}^{n}$ and $\mathsf{ESetInc}_{a,b,c,g}^{n}$ (or $\mathrm{GHD}_{a,b,a+b-2c,2g}$ and $\overline{\mathrm{GHD}}_{a,b,a+b-2c,2g}$ equivalently), the comparison between [23] and our results are shown as Table 1, where $n_{1},n_{2}$ are the smallest two numbers in ac,bc,c,nab+ca-c,b-c,c,n-a-b+c and $n_{1}\leq n_{2}$. It is worth noting that the results in [23] mainly depend on $n_{2}/g$, while our results mainly depend on $\sqrt{n_{1}n_{2}}/g$, which is a key factor in characterizing the quantum communication complexity of the Set-Inclusion Problem.
Table 1: The communication complexity of Set-Inclusion Problem
Communication complexity
Ref. [23] [^1] Ref. [23] gave the communication complexity of $\mathrm{GHD}_{a,b,c,g}$ and $\overline{\mathrm{GHD}}_{a,b,c,g}$ in Lemmas 3.3 and 3.4 actually. We reformulated their formulas for convenience. The details can be seen in Appendix D. Randomized: $O\left(\left(\frac{n_{2}}{g}\right)^{2}\log\frac{b}{g}\right)$ ($a\leq b\leq\frac{n}{2}$)
Randomized: $\Omega\left(\max\left\{\frac{n_{2}}{g},\log\left(\frac{\min\left\{a+b-2c,n-a-b+2c\right\}}{g}\right)\right\}\right)$
Our work Randomized: $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\right)$
Quantum: $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\right)$
Quantum: $\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)$
Remark 1.14.
It is worth noting that for both quantum and randomized communication complexities, the upper bound of $\mathsf{SetInc}_{a,b,c,g}^{n}$ is also the upper bound of $\mathsf{ESetInc}_{a,b,c,g}^{n}$. Similarly, the lower bound of $\mathsf{ESetInc}_{a,b,c,g}^{n}$ is also the lower bound of $\mathsf{SetInc}_{a,b,c,g}^{n}$.
@@ 2 Preliminaries
A multiset is a set with possibly repeating elements. We use $\{[\cdot]\}$ to denote multiset and $\{\cdot\}$ to denote standard set. Let $S$ be a multiset, $S\setminus\left\{a\right\}$ removes one occurrence of $a$ from $S$ if there is any.
@@ 2.1 Boolean Functions
A partial function is a function defined only on a subset of its domain $\mathcal{X}$. Formally, given a partial Boolean function $f\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$, $f(x)$ is undefined for $x\in\mathcal{X}$ if f(x)=f(x)=*. A total function is a function that is defined on the entire domain. We say $f\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$ is a subfunction of $g\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$ if $f(x)=g(x)$ or f(x)=f(x)=* for any $x\in\mathcal{X}$. For function $f\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$, we define $\overline{f}\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$ as
$\overline{f}(x)\coloneqq\begin{cases}-f(x)&\text{if $f(x)=1$ or $-1$},\\ *&\text{otherwise.}\end{cases}$
A Boolean predicate is a partial function that has domain $\mathcal{X}=\left\{0,1,...,n\right\}$ for any $n\in\mathbb{Z}^{+}$.
An incomplete Boolean matrix is a matrix with entries in $\left\{-1,1,*\right\}$, where undefined entries are filled with *.
@@
  1. 1.
    $\overline{x}\coloneqq(\overline{x_{0}},...,\overline{x_{n-1}})=(1-x_{0},...,1-x_{n-1})$;
  2. 2.
    $x\land y\coloneqq(x_{0}\land y_{0},...,x_{n-1}\land y_{n-1})$; and
  3. 3.
    $x\oplus y\coloneqq(x_{0}\oplus y_{0},...,x_{n-1}\oplus y_{n-1})$.
@@ 2.2 Communication Complexity Model
In the two-party communication model, Alice is given input $x\in\{0,1\}^{n}$, and Bob is given input $y\in\{0,1\}^{n}$. Then they aim to compute $f(x,y)$ for some function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ by communication protocols while minimizing the amount of communication between them. In this paper, we consider the communication protocols in the deterministic, randomized, and quantum settings, respectively. Furthermore, the formal definitions of communication complexities are shown as follows.
Definition 2.1 (Deterministic Communication Complexity).
For any function $f:\set{0,1}^{n}\times\set{0,1}^{n}\rightarrow\{-1,1,*\}$, the deterministic communication complexity $D(f)$ is defined as the cost of the deterministic protocol with the smallest cost, which computes $f$ correctly on any input. In particular,
$D(f)=\min_{\begin{subarray}{c}\Pi:\text{ deterministic protocol}\\ \forall x,y\in\{0,1\}^{n}\text{ s.t. }\Pi(x,y)=f(x,y)\end{subarray}}\mathrm{CC}\left(\Pi\right),$
where $\mathrm{CC}\left(\Pi\right)$ is the number of bits exchanged between two parties in the protocol $\Pi$.
Definition 2.2 (Randomized Communication Complexity).
For any function $f:\set{0,1}^{n}\times\set{0,1}^{n}\rightarrow\{-1,1,*\}$, the randomized communication complexity $R_{\epsilon}(f)$ is defined as the cost of the randomized protocol with the smallest cost, which has access to public randomness and computes f correctly on any input with probability at least $1-\epsilon$. In particular,
$R_{\epsilon}(f)=\min_{\begin{subarray}{c}\Pi:\text{ randomized protocol}\\ \forall x,y\in\{0,1\}^{n}\text{ s.t. }f(x,y)\neq*:\Pr[\Pi(x,y)]=f(x,y)]\geq 1-\epsilon\end{subarray}}\mathrm{CC}\left(\Pi\right),$
where $\mathrm{CC}\left(\Pi\right)$ is the number of bits exchanged between two parties in the protocol $\Pi$.
Definition 2.3 (Quantum Communication Complexity).
For any function $f:\set{0,1}^{n}\times\set{0,1}^{n}\rightarrow\{-1,1,*\}$, the quantum communication complexity $Q_{\epsilon}(f)$ is defined as the cost of the quantum protocol with the smallest cost, which is allowed to share prior entanglement and computes $f$ correctly on any input with probability at least $1-\epsilon$. In particular,
$Q_{\epsilon}(f)=\min_{\begin{subarray}{c}\Pi:\text{ quantum protocol}\\ \forall x,y\in\{0,1\}^{n}\text{ s.t. }f(x,y)\neq*:\Pr[\Pi(x,y)]=f(x,y)]\geq 1-\epsilon\end{subarray}}\mathrm{CC}\left(\Pi\right),$
where $\mathrm{CC}\left(\Pi\right)$ is the number of qubits exchanged between two parties in the protocol $\Pi$.
If a protocol succeeds with probability at least $1-\epsilon$ on any input for some constant $\epsilon<1/2$, we say the protocol is with bounded error. If $\epsilon=1/3$, we abbreviate $R_{\epsilon}(f),Q_{\epsilon}(f)$ as $R(f),Q(f)$. It is a folklore conclusion that $Q(f)\leq R(f)\leq D(f)$.
@@ 2.3 Permutation-Invariant Functions
By Remark 1.2, any permutation-invariant function $f:\left\{0,1\right\}^{n}\times\left\{0,1\right\}^{n}\rightarrow\left\{-1,1,*\right\}$ depends only on $|x|,|y|$ and $|x\land y|$. Thus, for any $a,b\in\left\{0,1,...,n\right\}$, there exists a function $f_{a,b}\colon\left\{\max\left\{0,a+b-n\right\},...,\min\left\{a,b\right\}\right\}\rightarrow\left\{-1,1{,*}\right\}$ such that
@@
Definition 2.4 (Jump in $f_{a,b}$).
For $f_{a,b}$ defined as (1), consider $c$ and $g$ such that c+g,cgc+g,c-g are in the domain of $f_{a,b}$, we say $(c,g)$ is a jump in $f_{a,b}$ if all the following three conditions are satisfied:
@@
$\mathscr{J}\left(f_{a,b}\right)\coloneqq\left\{(c,g):\begin{array}[]{c}f_{a,b}(c-g),f_{a,b}(c+g)\in\{0,1\}\\ f_{a,b}(c-g)\neq f_{a,b}(c+g)\\ \forall i\in(c-g,c+g),f_{a,b}(i)=*\end{array}\right\}.$
Remark 2.5.
If $(c,g)$ is a jump in $f_{a,b}$, then one of $\mathsf{ESetInc}_{a,b,c,g}^{n}$ and $\overline{\mathsf{ESetInc}}_{a,b,c,g}^{n}$ is a subfunction of $f$. Additionally, Definition 2.4 is adapted from [23], while [23] gives the definition of jumps in $h_{a,b}$, where $h_{a,b}(\Delta(x,y))=f(x,y)$ for permutaion-invariant function $f$ and $\Delta(x,y)$ is the Hamming distance between $x$ and $y$.
The following measure $m(\cdot)$ is used to capture the quantum communication complexity of permutation-invariant functions, which is inspired by the complexity measure introduced in [23].
Definition 2.6 (Measure $m(f)$).
Fix $n\in\mathbb{Z^{+}}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ be a permutation-invariant function. If $f$ is non-trivial, we define the measure $m(f)$ of $f$ as follows:
$m(f)\coloneqq\max_{\begin{subarray}{c}a,b\in\left\{1,2,...,n-1\right\}\\ (c,g)\in\mathscr{J}\left(f_{a,b}\right)\\ n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}\\ n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)\end{subarray}}\frac{\sqrt{n_{1}n_{2}}}{g}.$
If $f(x,y)$ only depends on $|x|$ and $|y|$, let $m(f)\coloneqq 0$.
Remark 2.7.
Definition 2.6 is motivated by the quantum lower bound of the Exact Set-Inclusion Problem (See Lemma 3.3). That is, $m(f)$ is the maximum of the quantum lower bound of all subfunctions of $f$ that are isomorphic to an Exact Set-Inclusion Problem. As a comparison, a similar measure in [23] is defined as the maximum of the randomized lower bound of all subfunctions of $f$ that are isomorphic to an Exact Gap-Hamming-Distance Problem.
@@ 3 Polynomial Equivalence on Communication Complexity of Permutation-Invariant Functions
To show the polynomial equivalence between quantum and randomized communication complexity of permutation-invariant functions as stated in Theorem 1.8, we prove the following two theorems (proved in Sections 3.1 and 3.2, respectively) for the quantum and randomized communication complexities of permutation-invariant functions using the measure in Definition 2.6.
@@
Fix $n\in{\mathbb{Z}^{+}}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ be a permutation-invariant function. We have
$Q(f)=\Omega(m(f)).$
@@
Fix $n\in{\mathbb{Z}^{+}}$. Given a permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ and the corresponding measure $m(f)$ defined in Definition 2.6, we have
  1. 1.
    $R(f)=O\left(m(f)^{2}\log^{2}n\log\log n+\log n\right)$, and
  2. 2.
    $Q(f)=O\left(m(f)\log^{2}n\log\log n+\log n\right)$.